# Uniform irreducibility of Galois action on the $\ell$-primary part of Abelian $3$-folds of Picard type

### v1

@@ 2.2. Shimura surfaces and families of abelian threefolds of Picard type
Taking quotient by the unique index two subgroup $\operatorname{PSL}_{2}(\mathbf{F}_{p})$ of $\operatorname{PGL}_{2}(\mathbf{F}_{p})$ yields a canonical homomorphism $\varepsilon_{A,p}:\operatorname{Gal}_{k}\to\{\pm 1\}$ and we let $\varepsilon_{A,D}=\prod_{p\mid D}\varepsilon_{A,p}:\operatorname{Gal}_{k}\to\{\pm 1\}$.
Note that, for any open compact subgroup $\tilde{K}$, $A$ defines a $k$-rational point in $Y_{\tilde{K}}$ if, and only if, the Galois representation on the adelic Tate module has image in $\tilde{K}$. Hence a point in $Y_{\widetilde{K}^{\prime\prime}}(k)$ corresponds precisely to an abelian $3$-fold AA over $k$ of Picard type having trivial $\varepsilon_{A,D}$.
@@
Let $k$ be a number field containing $M$ over which the connected component of $Y_{\widetilde{K}}$ are defined, and fix a connected component $Y$ of $Y_{\widetilde{K}}\times_{M}k$ and a smooth open $U$ of $Y$ endowed with an abelian scheme $f:A\to U$ of Picard type. Denote by $\eta$ the generic point of the smooth surface $U$. Fixing a closed geometric point $\bar{x}$ of $U$ the étale fundamental group sits in the middle of a short exact sequence
(9) $1\to\Pi_{1}(U_{\bar{k}},\bar{x})\to\Pi_{1}(U,\bar{x})\to\operatorname{Gal}_{k}=\Pi_{1}(\{x\},\bar{x})\to 1.$
The morphism $f:A\to U$ being proper and smooth, one can consider the étale sheaf $\mathrm{R}^{1}f_{*}\mathbf{Z}_{\ell}$ on $U$. As $U$ is geometrically connected we have $\Pi_{1}(U,\bar{x})\simeq\Pi_{1}(U,\bar{\eta})$ and the latter acts on
@@
Finally for any closed curve $C\subset U$ defined over $k$, there is a natural map $\Pi_{1}(C,\bar{x})\to\Pi_{1}(U,\bar{x})$ whose composition with $\rho_{U,\ell}$ is denoted $\rho_{C,\ell}$. As $f:A\to U$ is of Picard type, for any $x\in C(k)$
@@
By a series of results of Cadoret–Tamagawa (see [3, 4]), the set $C_{\rho}$ of all $x\in C(k)$ for which $\Gamma_{x}$ is not open in $\Gamma_{C}$ is finite and for all $x\in C(k)\setminus C_{\rho}$ the index $\Gamma_{C}:\Gamma_{x}]$ is uniformly bounded.
The Mumford–Tate group $\operatorname{MT}(A)$ of a polarized abelian variety $A$ over $\mathbf{C}$ is the smallest connected reductive subgroup of $\operatorname{GL}(\operatorname{H}_{1}(A,\mathbf{Q}))$ over $\mathbf{Q}$, whose $\mathbf{R}$-points contain the associated $\mathbf{R}$-morphism $h:\mathbf{C}^{\ast}\to\operatorname{GL}(\operatorname{H}_{1}(A(\mathbf{C}),\mathbf{R}))$ coming from the Hodge decomposition. If we assume further that $A$ is defined over a number field $k\subset\mathbf{C}$ finitely generated over $\mathbf{Q}$, then the image $\Gamma_{\ell}$ of $\operatorname{Gal}_{k}$ acting on $(T_{\ell}A)$ is an $\ell$-adic Lie group. By a theorem of Deligne [8, Chap. I.2], we have ${\rm Lie}(\Gamma_{\ell})_{\mathbf{Q}_{\ell}}\subset\operatorname{Lie}(\operatorname{MT}(A)\otimes_{\mathbf{Z}_{\ell}}\mathbf{Q}_{\ell})$ and the Mumford–Tate conjecture, known for abelian varieties of dimension at most $3$, asserts that they are equal (see e.g. [5]).
As the Mumford–Tate group of the (generic point of the) universal family $f:A\to U$ is given by $\widetilde{G}=\mathrm{GU}_{2,1}$, it follows from the above discussion that the Mumford–Tate group of any abelian $3$-fold of Picard type is a reductive subgroup of $\widetilde{G}$. We have the following trichotomy.
@@
Let $\mathfrak{g}$ be a reductive Lie subalgebra of $\mathfrak{gu}(3,k)$ defined over a characteristic 0 field $k$. If $\mathfrak{g}^{\prime}\subset\mathfrak{su}(3,k)$ denotes the semi-simple part of $\mathfrak{g}$, exactly one of the following holds:
@@
    $\mathfrak{g}^{\prime}=\{0\}$, i.e. $\mathfrak{g}$ is abelian,
@@
    $\mathfrak{g}^{\prime}$ is a form of $\mathfrak{sl}(2,k)$,
@@
    $\mathfrak{g}^{\prime}=\mathfrak{su}(3,k)$.
@@
If $\mathfrak{g}^{\prime}_{\mathbf{C}}=\{0\}$, then $\mathfrak{g}^{\prime}=\{0\}$, whereas if $\mathfrak{g}^{\prime}_{\mathbf{C}}=\mathfrak{sl}(3,\mathbf{C})$, then $\mathfrak{g}^{\prime}=\mathfrak{su}(3,k)$ for dimension reasons. In the remaining cases, using the well known fact that any proper non-zero semi-simple Lie subalgebra of $\mathfrak{sl}(3,\mathbf{C})$ is isomorphic to $\mathfrak{sl}(2,\mathbf{C})$, we deduce that $\mathfrak{g}^{\prime}$ is a form of $\mathfrak{sl}(2,k)$. ∎
If $A$ is (potentially) of CM type (resp. admits (potentially) a non-trivial CM quotient), then its Mumford–Tate group is of the first (resp. second) type. It is natural to ask whether Theorem B could be further refined for abelian $3$-folds without non-trivial CM quotients.
2.4. Galois stable lattices and rationality
Suppose henceforth that $k$ is a number field. Let $(A,\iota,\theta)$ be a polarized abelian $3$-fold of Picard type over $k$, and let $\eta:\mathbf{A}_{f}\otimes_{\mathbf{Q}}V\xrightarrow{\sim}V_{f}A$ be an isomorphism sending $\langle\cdot,\cdot\rangle_{A}$ to a $\mathbf{A}_{f}^{\times}$-multiple of $\langle\cdot,\cdot\rangle_{V}$. As $\operatorname{MT}(A,\iota,\theta)\subset\widetilde{G}$ by Deligne [8, Cor. 6.2], the action of $\operatorname{Gal}_{k}$ on the adelic Tate module $V_{f}A$, together with the choice of $\eta$, yields a continuous homomorphism:
$\rho_{A,f}:\operatorname{Gal}_{k}\longrightarrow\widetilde{G}(\mathbf{A}_{f}).$
[…]
Given any prime number $\ell$, the resulting continuous homomorphism:
$\rho_{\ell}=\rho_{A,\ell}:\operatorname{Gal}_{k}\longrightarrow\widetilde{G}(\mathbf{Q}_{\ell}),$
has compact image, which is thus necessarily contained in some maximal compact $\widetilde{K}_{\ell}$ of $\widetilde{G}(\mathbf{Q}_{\ell})$.
We will denote by $\overline{\rho}_{\ell}$ the composition of $\rho_{\ell}$ with the natural surjection of $\widetilde{K}_{\ell}$ onto its reductive quotient $\overline{G}_{\ell}$. Then $\overline{\rho}_{\ell}$ acts on $\mathcal{L}_{\ell}\otimes_{\mathbf{Z}_{\ell}}\mathbf{F}_{\ell}$, where $\mathcal{L}_{\ell}$ is a $\mathcal{O}\otimes\mathbf{Z}_{\ell}$-lattice whose stabilizer in $\widetilde{G}(\mathbf{Q}_{\ell})$ is $\widetilde{K}_{\ell}$. While in general $\overline{\rho}_{\ell}$ depends on the choice of $\widetilde{K}_{\ell}$, or equivalently on the choice of a $\operatorname{Gal}_{k}$-stable lattice $\mathcal{L}_{\ell}$, it follows from a Theorem of Brauer and Nesbitt that its semi-simplification is independent of these choices.
If $\ell$ splits in $M$ then $G(\mathbf{Q}_{\ell})=\operatorname{GL}_{3}(\mathbf{Q}_{\ell})$ and $\overline{\rho}_{\ell}:\operatorname{Gal}_{k}\longrightarrow\widetilde{\operatorname{GL}}_{3}(\mathbf{F}_{\ell})$ is absolutely irreducible if, and only if, there exists a unique, up to homothecy, $\rho_{\ell}(\operatorname{Gal}_{k})$-stable $\mathbf{Z}_{\ell}$-lattice $\mathcal{L}_{\ell}$.
If $\ell$ does not split in $M$, then $G(\mathbf{Q}_{\ell})$ has rank $1$ and every edge of the corresponding Bruhat-Tits tree links a vertex with reductive quotient $\overline{G}{}^{\circ}_{\ell}$ to a vertex with reductive quotient $\overline{G}{}^{\prime}_{\ell}$ (see §1.1). As $\rho_{\ell}(\operatorname{Gal}_{k})$ acts on the tree by isometries, if it fixes any two (or more) vertices, then it necessarily fixes an edge, hence its image in the reductive quotient of any fixed vertex would be reducible. Conversely, since no irreducible subgroup of $\overline{G}{}^{\circ}_{\ell}$ or of $\overline{G}{}^{\prime}_{\ell}$ does fix an adjacent vertices, one can characterize the representations $\rho_{\ell}$ fixing a unique vertex as follows.
Note that $\theta$ yields an integral pairing on $\operatorname{H}_{1}(A(\mathbf{C}),\mathbf{Z})$ inducing a pairing on $T_{\ell}A\simeq\operatorname{H}_{1}(A(\mathbf{C}),\mathbf{Z}_{\ell})$ for each $\ell$. If $\ell$ does not divide the degree of $\theta$, then $T_{\ell}A$ is self-dual, and one can chose $\widetilde{K}_{\ell}\simeq\widetilde{K}_{\ell}^{\circ}$.
Lemma 2.2.
Let $(A,\iota,\theta)$ be a polarized abelian $3$-fold of Picard type over a number field $k$.
  1. (i)
    $\overline{\rho}_{\ell}$ is absolutely irreducible if, and only if, the exists a unique, up to homothecy, $\rho_{\ell}(\operatorname{Gal}_{k})$-stable lattice. The latter is necessarily self-dual and $\widetilde{K}_{\ell}$ is conjugated to $\widetilde{K}_{\ell}^{\circ}$.
  2. (ii)
    Suppose $\ell$ be a prime that does not split in $M$. Then $\overline{\rho}_{\ell}(\operatorname{Gal}_{k})$ is an irreducible subgroup of $\overline{G}{}^{\prime}_{\ell}$ if, and only if, the exists a unique, up to homothecy, pair of $\rho_{\ell}(\operatorname{Gal}_{k})$-stable lattices. The latter are almost self-dual and $\widetilde{K}_{\ell}$ is conjugated to $\widetilde{K}^{\prime}_{\ell}$.
  3. (iii)
    If $\theta$ is principal, then one can chose $\widetilde{K}\simeq\widetilde{K}^{\circ}$, i.e., $(A,\iota,\theta)$ defines a $k$-rational point on $Y_{\widetilde{K}^{\circ}}$.
2.5. Lie images of Galois representations
The Lie algebra $\mathfrak{h}\subset\mathfrak{gu}(3,\mathbf{Z}_{\ell})$ of the image of $\rho_{A,\ell}$ is algebraic and $\mathfrak{h}_{\mathbf{Q}_{\ell}}=\mathbf{Q}_{\ell}\otimes_{\mathbf{Z}_{\ell}}\mathfrak{g}_{\mathbf{Z}_{\ell}}$ is reductive by Faltings [11, Thm. 3].
Serre [29, C.3.7] has defined an integral model of the Mumford–Tate group and refined Deligne’s theorem to show that its Lie algebra $\mathfrak{h}_{\mathbf{Z}_{\ell}}$ contains $\mathfrak{g}_{\mathbf{Z}_{\ell}}$ as a subgroup. The integral Mumford–Tate conjecture, known for abelian varieties of dimension $\leqslant 3$, asserts that the image is an open subgroup.
Theorem 2.3.
@@ 2.5. Lie images of Galois representations
Let $A$ be an abelian $3$-fold of Picard type defined over a field $k$. Then $\rho_{A,\ell}$ is potentially reducible if, and only if, $A$ has a non-trivial CM quotient.
@@
Suppose that $\mathfrak{g}_{\mathbf{Q}_{\ell}}$ is a proper subalgebra of $\mathfrak{gu}(3,\mathbf{Q}_{\ell})$. It is, by Proposition 2.1, either abelian or contains $\mathfrak{su}(3,\mathbf{Q}_{\ell})$ or its semisimple part is a form of $\mathfrak{sl}(2,\mathbf{Q}_{\ell})$. In the second case, as it contains homotheties by Bogomolov [2], it must contain $\mathfrak{su}(3,\mathbf{Q}_{\ell})\oplus\mathbf{Q}_{\ell}$. In each of the remaining cases, after extending scalars to $\overline{\mathbf{Q}}_{\ell}$, $\rho_{A,\ell}$ is potentially reducible, i.e., after possibly replacing $k$ by a finite extension, $\rho_{A,\ell}$ contains (as a direct factor by Faltings) a character $\chi_{\ell}:k^{\times}\backslash\mathbf{A}_{k}^{\times}\to\overline{\mathbf{Q}}_{\ell}^{\times}$. As a sub-representation of $\rho_{A,\ell}$, $\chi_{\ell}$ is unramified outside a finite set of places and its restriction decomposition groups at places above $\ell$ it Hodge-Tate with weights belonging to $\{0,-1\}$. In addition it is pure of weight 1-1. By Minkowski’s proof of the Dirichlet unit theorem, $\chi_{\ell}$ corresponds to an algebraic Hecke character $\chi:k^{\times}\backslash\mathbf{A}_{k}^{\times}\to\mathbf{C}^{\times}$ whose infinity component is necessarily of the form $\mathrm{N}_{\Phi^{\prime}}\circ\mathrm{N}_{k/L^{\prime}}$ where $\mathrm{N}_{\Phi^{\prime}}$ is the partial norm given by a CM type $\Phi^{\prime}$ for a CM field $L^{\prime}\subset k$. By [31, Lem. 2] replacing $(L^{\prime},\Phi^{\prime})$ by its double reflex yields the same infinite component, hence we may and do assume that $(L^{\prime},\Phi^{\prime})$ is a primitive, i.e., coincides with the reflex of a CM field $L$ endowed with a CM type $\Phi$. Further replacing $k$ by a finite (abelian) extension one can assume that $\chi_{f}$ takes values in $L^{\times}$. By Casselman (see [31, Thm. 6]), there exists an abelian variety $B$ defined over $k\supset L^{\prime}$ which is CM of type $(L,\Phi)$ and such that $\rho_{B,\ell}=\chi_{\ell}$, hence
$\operatorname{Hom}_{\operatorname{Gal}_{k}}(\rho_{A,\ell},\rho_{B,\ell})\neq\{0\}.$
@@
By Faltings one deduces that $\operatorname{Hom}_{k}(A,B)\neq\{0\}$, hence $A$ contains a non-trivial CM quotient $A^{\prime}$. One can assume that $A^{\prime}\neq A$, hence there exists an abelian variety $A^{\prime\prime}$ which is not of CM type and such that $A$ is isogenous to $A^{\prime}\times A^{\prime\prime}$, i.e., $V_{A,\ell}=V_{A^{\prime},\ell}\oplus V_{A^{\prime\prime},\ell}$. Furthermore since $\operatorname{Hom}_{k}(A^{\prime},A^{\prime\prime})=\{0\}$, one can show that the isogeny is $\mathcal{O}$-linear. Hence $A^{\prime\prime}$ admits multiplication by $\mathcal{O}$ and since it is not of CM type, it has dimension $2$, from which one deduces that $\mathfrak{h}_{\mathbf{Z}_{\ell}}=\mathfrak{g}(\mathfrak{u}(2)\times\mathfrak{u}(1),\mathbf{Z}_{\ell})$.
Then either $\rho_{A^{\prime\prime},\ell}$ is Lie surjective, in which case $\mathfrak{g}_{\mathbf{Z}_{\ell}}=\mathfrak{h}_{\mathbf{Z}_{\ell}}$, or else $\rho_{A^{\prime\prime},\ell}$ is reducible which by repeating the above argument would contradict $A^{\prime\prime}$ not being of CM type. ∎
Lemma 2.4.
Let $\widetilde{K}$ be an open compact subgroup of $\widetilde{G}(\mathbf{A}_{f})$ and let $x\in Y_{\widetilde{K}}(k)$ be such that $\operatorname{MT}(x)\neq\widetilde{G}$. Then $x$ belongs to a special subvariety defined over $k$.
Proof.
[…]
3. Irregularity for Picard modular surfaces
We denote by $q(X)$ the irregularity of a projective algebraic surface $X$ over $\mathbf{C}$, given by the dimension of $\operatorname{H}^{1}(X,\mathcal{O}_{X})$. If $X$ is smooth and projective then $q(X)=\dim\operatorname{H}^{0}(X,\Omega^{1}_{X})$.

### v2

@@ 2.1. Abelian $3$-folds and Shimura surfaces of Picard type
Taking quotient by the unique index two subgroup $\operatorname{PSL}(2,\mathbf{F}_{p})$ of $\operatorname{PGL}(2,\mathbf{F}_{p})$ yields a canonical homomorphism $\varepsilon_{A,p}:\operatorname{Gal}_{k}\to\{\pm 1\}$ and we let $\varepsilon_{A,D}=\prod_{p\mid D}\varepsilon_{A,p}:\operatorname{Gal}_{k}\to\{\pm 1\}$.
Using the observation made after (12), the points in $Y_{\widetilde{K}^{\prime\prime}}(k)$ corresponds precisely to an abelian $3$-fold AA over $k$ of Picard type having trivial $\varepsilon_{A,D}$.
2.2. Complex reflexions and families of abelian threefolds of Picard type
It is important to observe that even though for $\widetilde{K}\subset\widetilde{K}^{\circ}$ each point of $Y_{\widetilde{K}}(\mathbf{C})$ is associated to an abelian $3$-fold of Picard type, there does not exist such a family over the entire $Y_{\widetilde{K}}(\mathbf{C})$ unless there are no points with extra automorphisms, in which case $Y_{\widetilde{K}}(\mathbf{C})$ would be a fine moduli space. In our cases of interest $\widetilde{K}$ is not neat, and therefore $Y_{\widetilde{K}}$ is not a fine moduli space. In this subsection we prove that there exists a natural family of abelian $3$-folds of Picard type over $Y_{\widetilde{K}^{\prime\prime}}$ minus a finite number of points (see (13)). This will be crucially used in the proof of Theorem B.
Recall that an element $\gamma$ of the discrete subgroup $\Gamma=\widetilde{G}(\mathbf{Q})\cap\widetilde{K}\widetilde{G}(\mathbf{R})$ has a fixed point in $\mathcal{H}$ if, and only if, $\gamma$ has finite order (this is because the stabilizers in $\widetilde{G}(\mathbf{R})$ of points in $\mathcal{H}$ are maximal compact subgroups). Such a $\gamma$ is called elliptic if it only fixes an isolated point in $\mathcal{H}$, otherwise it is called a complex reflexion. By [17, Cor. 4.5.10], the set of singular points of $Y_{\Gamma}$ is finite and consists of isolated fixed points of elliptic elements of $\Gamma$, all defined over a finite extension of $M$. Moreover, smooth points at which universal cover $\mathcal{H}\to Y_{\Gamma}$ is not étale are fixed points of complex reflexions, i.e.  order $2$ elements in $\Gamma$ fixing hyperbolic planes (see [17, §4.5]).
[…]
Lemma 2.1.
Proof.
Consider a complex reflexion $\gamma\in\Gamma^{\prime\prime}$ as an endomorphism of the Hermitian space $M^{3}$ having signature $(2,1)$. As the eigenspaces of $\gamma$ are mutually orthogonal, at most one such eigenspace could contain a negative line (corresponding to a point in $\mathcal{H}$). This if $\gamma$ fixes more that one point of $\mathcal{H}$, it necessarily fixes a hyperbolic line in $\mathcal{H}$. The corresponding endomorphism of $M^{3}$ has a mutually orthogonal eigenplane and eigenline (both $M$-rational), forcing the eigenvalues to be in $\mathcal{O}_{M}^{\times}=\{\pm 1\}$ (as $D>4$) and not all equal. It follows, that for any $p\mid D$, the image of $g_{f}^{-1}\gamma g_{f}\in\widetilde{K}^{\circ}$ into the projectivization of the reductive quotient of $\widetilde{K}_{p}^{\circ}$ belongs to the image under the adjoint isomorphism $\operatorname{PGL}(2,\mathbf{F}_{p})\xrightarrow{\sim}\mathrm{PGO}(3,\mathbf{F}_{p})$ of an element represented by a matrix having both eigenvalues $1$ and 1-1. In particular its image in $\operatorname{PGL}(2,\mathbf{F}_{p})/\operatorname{PSL}(2,\mathbf{F}_{p})=\{\pm 1\}$ equals $\left(\tfrac{-1}{p}\right)$. As $\prod_{p\mid D}\left(\tfrac{-1}{p}\right)=\left(\tfrac{-1}{D}\right)=-1$ it follows from (13) that $g_{f}^{-1}\gamma g_{f}\notin\widetilde{K}^{\prime\prime}$, i.e., $\gamma\notin\Gamma^{\prime\prime}$. ∎
Proposition 2.2.
There exists a finite extension $k$ of $M$ and a natural family of abelian $3$-folds of Picard type over $Y_{\widetilde{K}^{\prime\prime}}\times_{M}k$ minus a finite number of $k$-rational elliptic points.
Proof.
We claim that there is a family of abelian $3$-folds of Picard type over any open subset $U$ of $Y_{\widetilde{K}^{\prime\prime}}$ which contains no point with a non-trivial stabilizer. By [18, §2.3.4], the moduli stack $S_{\widetilde{K}^{\prime\prime}}$ associated to this problem is an algebraic stack (for the étale topology), locally of finite type over the base which we may take to $\operatorname{Spec}(M)$. Moreover, by [18, §A.7.5], there is a canonical surjective morphism $\phi$ from $S_{\widetilde{K}^{\prime\prime}}$ to the associated coarse moduli space $S_{\widetilde{K}^{\prime\prime}}$], which in our notations is $Y_{\widetilde{K}^{\prime\prime}}$. By [18, §7], $S_{\widetilde{K}^{\prime\prime}}$] is an algebraic space and even a quasi-projective scheme. Moreover, by a general property of moduli stacks (see [23, Chap. 7]), $\phi$ is an isomorphism over the locus $U$ where there is no non-trivial automorphism, by which we mean it has no infinitesimal automorphism; analytically, this corresponds to points of $Y_{\widetilde{K}}(\mathbf{C})$ having no non-trivial stabilizers. Now $U$ is a priori an open subscheme of $S_{\widetilde{K}^{\prime\prime}}$], but since it is where $\phi$ is an isomorphism, we get a canonical open $U\hookrightarrow S_{\widetilde{K}^{\prime\prime}}$. This map tautologically yields the desired family $f:A\to U$ of abelian varieties of Picard type. In view of Lemma 2.1 and the discussion preceding it, after possibly enlarging $k$, we may consider $U$ be the complementary open in $Y^{\prime\prime}$ of its finitely many elliptic points. ∎
@@ 2.2. Complex reflexions and families of abelian threefolds of Picard type
Let $k$ be a number field containing $M$ over which the connected component of $Y_{\widetilde{K}}$ are defined. Fix a connected component $Y$ of $Y_{\widetilde{K}}\times_{M}k$ and a smooth open $U$ of $Y$ endowed with an abelian scheme $f:A\to U$ of Picard type. Denote by $\eta$ the generic point of the smooth surface $U$. Fixing a closed geometric point $\bar{x}$ of $U$ the étale fundamental group sits in the middle of a short exact sequence
(16) $1\to\Pi_{1}(U_{\bar{k}},\bar{x})\to\Pi_{1}(U,\bar{x})\to\operatorname{Gal}_{k}\to 1.$
The morphism $f:A\to U$ being proper and smooth, one can consider the étale sheaf $\mathrm{R}^{1}f_{*}\mathbf{Z}_{\ell}$ on $U$. As $U$ is geometrically connected, we have $\Pi_{1}(U,\bar{x})\simeq\Pi_{1}(U,\bar{\eta})$ and the latter acts on
@@
Finally for any closed curve $C\subset U$ defined over $k$, there is a natural map $\Pi_{1}(C,\bar{x})\to\Pi_{1}(U,\bar{x})$ whose composition with $\rho_{U,\ell}$ is denoted $\rho_{C,\ell}$. As $f:A\to U$ is of Picard type, for any $x\in C(k)$,
@@
By a series of results of Cadoret–Tamagawa (see [3, 4]), the set $C_{\rho}$ of all $x\in C(k)$ for which $\Gamma_{x}$ is not open in $\Gamma_{C}$ is finite and the index $\Gamma_{C}:\Gamma_{x}]$ is uniformly bounded for $x\in C(k){\smallsetminus}C_{\rho}$.
Recall that the Mumford–Tate group $\operatorname{MT}(A)$ of a polarized abelian variety $A$ over $\mathbf{C}$ is the smallest connected reductive subgroup of $\operatorname{GL}(\operatorname{H}_{1}(A,\mathbf{Q}))$ over $\mathbf{Q}$, whose $\mathbf{R}$-points contain the image of the associated $\mathbf{R}$-morphism $h:\mathbf{C}^{\times}\to\operatorname{GL}(\operatorname{H}_{1}(A(\mathbf{C}),\mathbf{R}))$ coming from the Hodge decomposition. If $A$ is defined over a field $k$ finitely generated over $\mathbf{Q}$, then the image $\Gamma_{\ell}$ of $\operatorname{Gal}_{k}$ acting on $T_{\ell}A$ is an $\ell$-adic Lie group. Denoting $\mathfrak{g}_{\mathbf{Z}_{\ell}}$ its Lie algebra, it is a theorem of Deligne [9, Chap. I.2] that $\mathfrak{g}_{\mathbf{Q}_{\ell}}=\mathfrak{g}_{\mathbf{Z}_{\ell}}\otimes_{\mathbf{Z}_{\ell}}\mathbf{Q}_{\ell}\subset\operatorname{Lie}(\operatorname{MT}(A)\otimes_{\mathbf{Q}}\mathbf{Q}_{\ell})$ and the Mumford–Tate conjecture, known for abelian varieties of dimension at most $3$, asserts that they are equal (see e.g. [5]).
As the Mumford–Tate group of the (generic point of the) universal family $f:A\to U$ is given by $\widetilde{G}=\mathrm{GU}(2,1)$, it follows from Deligne [9, Cor. 6.2] that the Mumford–Tate group of any abelian $3$-fold of Picard type is a reductive subgroup of $\widetilde{G}$. We have the following trichotomy.
@@
Let $\mathfrak{g}$ be a reductive Lie subalgebra of $\mathfrak{gu}(3,\mathbf{Q}_{\ell})$. If $\mathfrak{g}^{\prime}\subset\mathfrak{su}(3,\mathbf{Q}_{\ell})$ denotes the semi-simple part of $\mathfrak{g}$, exactly one of the following holds:
@@
    $\mathfrak{g}^{\prime}=\{0\}$, i.e., $\mathfrak{g}$ is abelian,
@@
    $\mathfrak{g}^{\prime}$ is a form of $\mathfrak{sl}(2,\mathbf{Q}_{\ell})$,
@@
    $\mathfrak{g}^{\prime}=\mathfrak{su}(3,\mathbf{Q}_{\ell})$.
@@
If $\mathfrak{g}^{\prime}_{\overline{\mathbf{Q}}_{\ell}}=\{0\}$, then $\mathfrak{g}^{\prime}=\{0\}$, whereas if $\mathfrak{g}^{\prime}_{\overline{\mathbf{Q}}_{\ell}}=\mathfrak{sl}(3,\overline{\mathbf{Q}}_{\ell})$, then $\mathfrak{g}^{\prime}=\mathfrak{su}(3,\mathbf{Q}_{\ell})$ for dimension reasons. In the remaining cases, using the well known fact that any proper non-zero semi-simple Lie subalgebra of $\mathfrak{sl}(3,\overline{\mathbf{Q}}_{\ell})$ is isomorphic to $\mathfrak{sl}(2,\overline{\mathbf{Q}}_{\ell})$, we deduce that $\mathfrak{g}^{\prime}$ is a form of $\mathfrak{sl}(2,\mathbf{Q}_{\ell})$. ∎
If $A$ is of CM type (resp. admits a CM factor), then $\operatorname{Lie}(\operatorname{MT}(A)\otimes_{\mathbf{Q}}\mathbf{Q}_{\ell})$ is of the first (resp. second) type in Lemma 2.3. Also clearly $\rho_{A,\ell}$ is potentially abelian (resp. is potentially reducible) if, and only if, $\mathfrak{g}_{\mathbf{Q}_{\ell}}$ is of the first (resp. second) type in Lemma 2.3. The following proposition bridges the two sides.
Proposition 2.4.
@@
Let $A$ be an abelian $3$-fold of Picard type over a number field $k$. Then $\rho_{A,\ell}$ is potentially abelian (resp. is potentially reducible) if, and only if, $A$ is of CM type (resp. admits a CM factor).
@@
One implication is clear. For the other, after extending scalars to $\overline{\mathbf{Q}}_{\ell}$ and after possibly replacing $k$ by a finite extension, we may assume that $\rho_{A,\ell}$ is potentially reducible, i.e. it contains (as a direct factor by Faltings) a character $\chi_{\ell}:k^{\times}\backslash\mathbf{A}_{k}^{\times}\to\overline{\mathbf{Q}}_{\ell}^{\times}$. Being a sub-representation of $\rho_{A,\ell}$, $\chi_{\ell}$ is unramified outside a finite set of places, and its restriction to decomposition groups at places above $\ell$ is Hodge-Tate with weights 0 and 1-1. In addition it is pure of weight 1-1. By Weil, $\chi_{\ell}$ corresponds to an algebraic Hecke character $\chi:k^{\times}\backslash\mathbf{A}_{k}^{\times}\to\mathbf{C}^{\times}$ whose infinity component is necessarily of the form $\mathrm{N}_{\Phi^{\prime}}\circ\mathrm{N}_{k/L^{\prime}}$ where $\mathrm{N}_{\Phi^{\prime}}$ is the partial norm given by a CM type $\Phi^{\prime}$ for a CM field $L^{\prime}\subset k$. By [31, Lem. 2] replacing $(L^{\prime},\Phi^{\prime})$ by its double reflex yields the same infinite component, hence we may and do assume that $(L^{\prime},\Phi^{\prime})$ is primitive, i.e., coincides with the reflex of a CM field $L$ endowed with a CM type $\Phi$. Further replacing $k$ by a finite (abelian) extension one can assume that $\chi_{f}$ takes values in $L^{\times}$. By Casselman (see [31, Thm. 6]), there exists an abelian variety $A^{\prime}$ defined over $k$ which is CM of type $(L,\Phi)$ and such that $\rho_{A^{\prime},\ell}=\chi_{\ell}$, hence
$\operatorname{Hom}_{\operatorname{Gal}_{k}}(\rho_{A,\ell},\rho_{A^{\prime},\ell})\neq\{0\}.$
@@
By Faltings, one deduces that $\operatorname{Hom}_{k}(A,A^{\prime})\neq\{0\}$, hence $A$ contains a non-trivial CM quotient. If $A$ is of CM type we’re done. If not, then there exists an abelian variety $A^{\prime\prime}$ which is not of CM type and such that $A$ is isogenous to $A^{\prime}\times A^{\prime\prime}$, i.e., $V_{\ell}A=V_{\ell}A^{\prime}\oplus V_{\ell}A^{\prime\prime}$. Furthermore since $\operatorname{Hom}_{k}(A^{\prime},A^{\prime\prime})=\{0\}$, one can show that the isogeny is $\mathcal{O}_{M}$-linear. Hence $A^{\prime\prime}$ is also of Picard type, from which one deduces that we’re in the second case in Lemma 2.3, in particular $\rho_{A,\ell}$ is not potentially abelian. ∎
3. Uniform irreducibility of Galois images